# LSAQ: Layer-Specific Adaptive Quantization for Large Language Model Deployment

### v1

@@ III-B Resource Detection Module
Effective management of GPU resources is critical for the edge-side deployment of LLMs, as it directly impacts the feasibility of deployment as well as the subsequent performance and responsiveness of the models. The Resource Detection (RD) module is designed to adapt to GPU devices of various scales and configurations, whether deployed on personal computers, servers, or cloud computing environments.
Through flexible environment variable configurations and programmatic approaches, the resource detection module can automatically select the GPU device with the most available free memory. This capability ensures efficient loading and inference of quantized LLMs, optimizing their performance during deployment.
TABLE I: Zero-shot QA task results of quantized Llama models
@@
III-C Quantization Strategy Formulation Module
The core of the Quantization Strategy Formulation (QSF) module lies in designing an optimal model quantization deployment plan based on the layer importance of the target LLM and the current GPU resource availability. This strategy aims to ensure that the quantized model minimizes accuracy loss while enabling seamless deployment and efficient inference execution.
Specifically, this module first evaluates the relationship between GPU memory availability and the memory requirements of the model during loading and inference. If the GPU memory is sufficient, the model is deployed without quantization to maintain optimal performance. If the memory is insufficient for an unquantized model but sufficient for an INT8-quantized model, the entire model undergoes INT8 quantization, as the accuracy loss from INT8 quantization is negligible.
Further, if the GPU memory is insufficient for an INT8-quantized model but can support an INT4-quantized model, the module applies INT4 quantization to layers with low importance while retaining INT8 quantization for the remaining layers. Since INT4 quantization has a greater impact on model accuracy, the strategy aims to maximize the proportion of INT8-quantized layers while ensuring that the quantized model’s memory requirements remain within the GPU’s capacity. Finally, if the GPU memory cannot accommodate even an INT4-quantized model, the quantization process will be paused until other processes release memory.
This module aims to achieve an optimal balance between model accuracy and hardware resources, ensuring adaptability to various deployment environments.
III-D Model Quantization Module
After determining the deployment strategy via the QSF module, this module performs per-channel quantization on the layers of LLMs that require quantization. Compared to the per-tensor quantization method, per-channel quantization assigns independent quantization parameters to each channel of the weight matrix. This approach allows the quantization process to more accurately capture the dynamic range of data in each channel, thereby reducing quantization errors and improving model accuracy. The quantization process can be mathematically represented as:
where $W$ denotes the weight matrix of the model, $W_{i}$ represents all values in the $i$-th row of the weight matrix, and the superscript of $W$ specifies its associated data type.
IV Experiments
@@ IV-A Experimental Setup
IV-A1 Models
To verify the feasibility of our method, we conducted experiments on some of the most popular open-source LLMs in the field of natural language processing, including Llama2-7B, Llama2-13B, and Llama3-8B, developed by Meta AI. These LLMs are based on the decoder-only Transformer architecture. Due to their extensive pretraining on large-scale language data, they demonstrate exceptional performance across a variety of NLP tasks.
@@ IV-A2 Benchmarks
To comprehensively evaluate the performance of LLMs after quantization, we adopted a dual evaluation strategy.
Firstly, we selected six zero-shot tasks to evaluate whether the core capabilities of the models are preserved after quantization:
  • PIQA [25]: A physical interaction question-answering dataset focuses on testing the model’s ability to predict how objects interact in the physical world. The questions are typically based on everyday physical scenarios, such as the typical uses of objects or non-typical but practically feasible uses. PIQA consists of approximately 20,000 question-answer pairs, which can be multiple-choice or true/false questions.
@@
  • BoolQ [27]: A question-answering task where the model must predict whether a passage contains the answer to a specific question. These questions are typically complex and non-factual queries, requiring a ”yes” or ”no” response based on the given passage.
@@
  • WinoGrande [29]: An open-domain language understanding dataset consisting of 1.5 million natural language sentence fragments, each ranging from 4 to 14 words in length. These fragments are annotated as ”yes” or ”no” to indicate whether they align with commonsense or world knowledge.
@@
Additionally, we have assessed the perplexity (PPL) of quantized LLMs on the WikiText2 [30] dataset. This dataset, introduced by Merity et al., is derived from Wikipedia articles and encompasses a wealth of encyclopedic knowledge. By measuring the PPL of LLMs in language generation tasks across various levels of quantization precision, we are able to accurately evaluate the models’ predictive capabilities on the test data.
@@ IV-A3 Baseline
In the experiment, we selected Layer-Wise Quantization (hereinafter referred to as LWQ) as the baseline method. The LWQ technique aligns with our quantization strategy in terms of granularity, as both apply differential quantization based on the importance of each layer. The LWQ method evaluates layer importance using two different metrics: one is the cosine similarity between the layer’s input and output, and the other is the number of weights in the layer that are significantly greater than the average value.
Experimental results from the LWQ paper indicate that using cosine similarity as the basis for quantization leads to better performance for the Llama series models. Based on this consideration, we decided to use cosine similarity as the layer importance metric for LWQ in formulating our quantization strategy. Since this method is independent of the underlying quantization technique, we will test it using the same quantization approach as ours.
@@ IV-B Main Results
Tables I and III present the detailed results of our experiments. To comprehensively evaluate the application of LSAQ quantization technology to LLMs, we tested the zero-shot task accuracy and perplexity of quantized Llama-2 (7B/13B) and Llama-3 (8B) models. In these experiments, INT4 quantization was applied to 25%, 50%, and 75% of the model’s layers, while the remaining layers used INT8 quantization. This quantization scheme reduced the average bit-width of LLMs to 7 bits, 6 bits, and 5 bits, corresponding to different quantization ratios.
@@ IV-B1 Results on Zero-shot tasks
Table I provides a detailed breakdown of the accuracy and overall average scores of various quantized models across multiple zero-shot question-answering tasks.
However, there is a special case in the experimental results. On the Llama-2-7B model, the performance of the two quantization methods is exactly the same under the condition of 7-bit quantization. This is because, although there are differences in determining the order of the first 25% least important layers between the method using cosine similarity of inputs and outputs and the method using Jaccard similarity with the vocabulary, they ultimately reach a consensus at the overall level, as shown in Table II. Consequently, both methods exhibit a high degree of consistency across all performance metrics. This consistency is reflected in both the zero-shot evaluation experiments and the subsequent perplexity experiments.
TABLE II: The First 25% Layer Numbers of Llama-2-7B under Different Layer Importance Indicators
@@
From the experimental data, it can be observed that the LSAQ (our) quantized model generally achieves superior performance in these zero-shot tasks. Except for the 25% quantization case of Llama-2-7B (where both methods show high consistency), the LSAQ method outperforms the LWQ method in terms of average accuracy across all tasks and has an advantage in accuracy for 87.5% of individual tasks. Notably, for the ARC-e and HellaSwag tasks, our method achieves higher accuracy under a wide range of conditions. Moreover, the LSAQ method is able to maintain a high level of accuracy even at lower bit numbers.
@@ IV-B2 Results on Perplexity
The PPL of a model is a critical metric in the field of NLP used to evaluate the performance of language models. It measures the model’s ability to predict a set of sample data, particularly in assessing the model’s predictive accuracy and generalization capability. A lower perplexity value indicates higher prediction accuracy, thereby reflecting the impact of quantization on model performance.
Table III demonstrates the impact of different quantization methods on the perplexity of three LLMs at various quantization bit widths. The experimental results indicate that, except for the condition where Llama-3-8B is quantized to 7 bits, where the LWQ quantization slightly outperforms LSAQ, in the remaining 90% of cases, the model quantized by LSAQ exhibits lower perplexity on the Wikitext2 dataset.
TABLE III: Perplexity Results of Quantized Llama Models on the Wikitext2 Dataset
@@ IV-B3 Analysis of results
By analyzing the experimental data of the quantized model on zero-shot tasks and model perplexity, it can be demonstrated that evaluating the layer importance of LLMs by constructing the top-k token sets corresponding to the input and output of each layer and calculating the Jaccard similarity of these sets is effective. This evaluation approach adeptly captures the interplay between each layer and semantic information. Consequently, it facilitates a more nuanced and accurate appraisal of the significance of each layer.
This finding reveals that compared to LWQ, the LSAQ quantization method is more effective in maintaining model accuracy, as it loses less critical information during the quantization process. Consequently, this suggests that LSAQ has potential advantages in the field of model compression and deployment on edge devices.
IV-C Quantized Model Deployment
In addition to quantization, dynamic deployment of LLMs on edge devices is also one of the core capabilities of LSAQ. To investigate whether our system meets the criteria for successful deployment of large models on edge devices, we conducted tests on the memory requirements for loading model weights of the quantized Llama-2 (7B/13B) and Llama-3 (8B) models. Besides, we examined the quantization strategies provided by the system for the Llama-2-7B model under various memory constraints to assess whether they are compliant with the requirements.
@@ IV-C Quantized Model Deployment
Figure 4 clearly demonstrates the amount of memory required to load the weights of three models at different quantization precisions. When these models operate at FP16 precision, they require approximately 12.82GB, 24.36GB, and 15.14GB of memory, respectively. With the reduction of the average number of bits after quantization, the memory occupancy of all models shows a significant decrease. At their lowest, the memory requirements can be reduced to about 3.56GB, 6.79GB, and 5.18GB, respectively, which allows these models to be easily deployed on mainstream graphics cards currently available on the market.
TABLE IV: Quantization Strategies for Llama-2-7B under Varying Memory Conditions
@@
During the inference process of LLMs, intermediate computational results may be generated, which also occupy a certain amount of memory. Consequently, when devising our quantization strategy, we took into full consideration the additional demand for memory during the inference process of LLMs. Consequently, we have designed corresponding safety memory strategies based on the requirements of models of different scales.
Table IV illustrates the actual quantization strategies employed by LSAQ for deploying Llama-2-7B under various memory configurations. For instance, if a GPU has 16GB of available memory, it is entirely feasible to deploy Llama-2-7B directly on the GPU at FP16 precision without the need for quantization. However, when the GPU has only 6GB of available memory, it becomes necessary to quantize the least important 10 out of the 32 layers of Llama-2-7B to INT4 precision, while the remaining 22 layers are quantized to INT8 precision, thereby facilitating the deployment of the model.
@@ V Conclusion
This paper introduces LSAQ, a layer-specific adaptive quantization and dynamic deployment system for large language models. LSAQ adjusts quantization strategies on hardware platforms with varying resource sizes by establishing a novel layer importance assessment mechanism, thereby significantly reducing the model’s storage requirements while maintaining model performance as much as possible. Experimental results demonstrate that the accuracy of the LSAQ-quantized model on zero-shot tasks and the perplexity of the language model are superior to existing quantization methods of the same granularit, even at lower bit widths, thus maintaining high performance. This indicates that LSAQ has potential advantages in model compression and on-device model deployment. By employing intelligent adaptive quantization methods, LSAQ can efficiently deploy quantized models in resource-constrained environments, providing a new perspective and approach for the deployment of large language models on edge devices.

### v2

@@ III-B Resource Detection Module
Effective management of GPU resources is critical for the edge-side deployment of LLMs, as it directly impacts the feasibility of deployment as well as the subsequent performance and responsiveness of the LLMs. The Resource Detection (RD) module is designed to adapt to GPU devices of various scales and configurations, including scenarios of personal computers, servers, or cloud computing environments.
Through flexible environment variable configurations and programmatic approaches, the RD module can automatically select the GPU device with the most available free memory. This capability ensures efficient loading and inference of quantized LLMs, optimizing their performance during deployment.
III-C Quantization Strategy Formulation Module
The core of the Quantization Strategy Formulation (QSF) module lies in designing an optimal LLM quantization deployment plan based on the layer importance of the target LLM and the current GPU resource availability. This strategy aims to ensure that the quantized LLM minimizes accuracy loss while enabling seamless deployment and efficient inference execution. We present the complete pseudocode for QSF in Algorithm 1.
Specifically, the module first assesses whether the LLM can be deployed directly without quantization or after quantization to INT8, based on the available GPU resources. In both deployment scenarios, the accuracy loss of the LLM is negligible. Furthermore, if the GPU resources are insufficient to support the two aforementioned types of precision models but can accommodate the INT4 quantization model, the module will apply INT4 quantization to the less important layers. The remaining layers will continue to use the INT8 quantization. This approach maximizes the proportion of INT8 quantized layers to maintain the accuracy of the LLM. Finally, if the GPU resources cannot meet any of the above conditions, the quantization task will be paused and wait for other processes to release the necessary resources.
This module aims to achieve an optimal balance between LLM accuracy and GPU resources, ensuring adaptability to various deployment environments.
III-D Model Quantization Module
After determining the deployment strategy via QSF, this module performs per-channel quantization on the layers of LLMs that require quantization. Compared to the per-tensor quantization method, per-channel quantization assigns independent quantization parameters to each channel of the weight matrix. This approach allows the quantization process to more accurately capture the dynamic range of data in each channel, thereby reducing quantization errors and improving LLM accuracy. The quantization process can be mathematically represented as:
where $W$ denotes the weight matrix of the LLM, $W_{i}$ represents all values in the $i$-th row of the weight matrix, and the superscript of $W$ specifies its associated data type.
IV Experiments
TABLE I: Zero-shot QA task results of quantized Llama models. $\uparrow$ denotes the larger the better. Bold values denote the best performance scores.
@@ IV-A Experimental Setup
IV-A1 LLMs
To verify the feasibility of our method, we conducted experiments on state-of-the-art open-source LLMs, including Llama-2-7B, Llama-2-13B, and Llama-3-8B, developed by Meta AI. These LLMs are based on the decoder-only Transformer architecture. Due to their extensive pre-training on large-scale language data, they demonstrate exceptional performance across a variety of natural language processing tasks.
@@ IV-A2 Benchmarks
To comprehensively evaluate the performance of LLMs after quantization, we adopted a dual evaluation strategy. Firstly, we selected six zero-shot tasks to evaluate whether the core capabilities of the LLMs are preserved after quantization:
  • PIQA [28]: A physical commonsense reasoning dataset that tests a model’s ability to choose appropriate solutions in everyday physical scenarios by providing a physical goal and two possible solutions. It emphasizes understanding the properties and manipulation of objects and contains over 16,000 question-answer pairs to assess the model’s physical commonsense reasoning ability.
@@
  • BoolQ [30]: A question-answering task where the model must predict whether a passage contains the answer to a specific question. These questions require a ”yes” or ”no” response based on the given passage.
@@
  • WinoGrande [32]: A large-scale commonsense reasoning dataset, based on the design of the Winograd Schema Challenge, which uses crowdsourcing and bias reduction algorithms to generate 44,000 problems. It aims to more accurately assess the common sense reasoning capabilities of artificial intelligence by increasing the scale and difficulty of the problems.
@@
Secondly, we have assessed the perplexity (PPL) of quantized LLMs on the WikiText2 [20] dataset. This dataset is derived from Wikipedia articles and encompasses a wealth of encyclopedic knowledge. By measuring the PPL of LLMs in language generation tasks across various levels of quantization precision, we are able to accurately evaluate the LLMs’ predictive capabilities on the test data.
@@ IV-A3 Baseline
In the experiment, we selected Layer-Wise Quantization (hereinafter referred to as LWQ) as the baseline method. The LWQ technique aligns with our quantization strategy in terms of granularity, since they both apply differential quantization based on the importance of each layer. The LWQ method evaluates layer importance using two different metrics: one is the cosine similarity between the layer’s input and output, and the other is the number of weights in the layer that are significantly greater than the average value.
Experimental results from the LWQ paper indicate that using cosine similarity as the basis for quantization leads to better performance for the Llama series models. On the basis of this consideration, we decided to use cosine similarity as the layer importance metric for LWQ in formulating a quantization strategy. Since this method is independent of the underlying quantization technique, we will test it using the same quantization approach as ours.
@@ IV-B Main Results
Tables I and III present the detailed results of our experiments. To comprehensively evaluate the application of LSAQ quantization technology to LLMs, we tested the zero-shot task accuracy and perplexity of quantized Llama-2-7B/13B and Llama-3-8B models. In these experiments, INT4 quantization was applied to 25%, 50%, and 75% of LLMs’ layers, while the remaining layers used INT8 quantization. This quantization scheme reduced the average bit-width of LLMs to 7 bits, 6 bits, and 5 bits, corresponding to different quantization ratios.
@@ IV-B1 Results on zero-shot tasks
Table I provides a detailed breakdown of the accuracy and overall average scores of various quantized LLMs across multiple zero-shot question-answering tasks.
TABLE II: The first 25% layer numbers of Llama-2-7B under different layer importance indicators.
@@
However, there is a special case in the experimental results. On the Llama-2-7B model, the performance of the two quantization methods is exactly the same under the condition of 7-bit quantization. We attribute it to the fact that although there are differences in determining the order of the first 25% least important layers with two indicators, they ultimately reach a consensus at the overall level, as shown in Table II. Consequently, both methods exhibit a high degree of consistency across all performance metrics. This consistency is reflected in both the zero-shot evaluation experiments and the subsequent perplexity experiments.
From the experimental data, it can be observed that the LSAQ quantized LLM generally achieves superior performance in these zero-shot tasks. Except for the 25% quantization case of Llama-2-7B (where both methods show high consistency), the LSAQ method outperforms the LWQ method in terms of average accuracy across all tasks and has an advantage in accuracy for 87.5% of individual tasks. Notably, for the ARC-e and HellaSwag tasks, our method achieves higher accuracy under a wide range of conditions. Moreover, the LSAQ method is able to maintain a high level of accuracy even at lower bit numbers.
@@ IV-B2 Results on perplexity
The PPL is a critical metric in the field of NLP used to evaluate the performance of LLMs. It measures the LLM’s ability to predict a set of sample data, particularly in assessing the LLM’s predictive accuracy and generalization capability. A lower perplexity value indicates higher prediction accuracy, thereby reflecting the impact of quantization on LLM performance.
Table III demonstrates the impact of different quantization methods on the perplexity of three LLMs at various quantization bit widths. The experimental results indicate that, except for the condition of $<$Llama-3-8B, 7 bits$>$ where the LWQ quantization slightly outperforms LSAQ, the LLM quantized by LSAQ exhibits lower perplexity on the WikiText2 dataset in all the remaining cases.
TABLE III: Perplexity results of quantized Llama models on the WikiText2 dataset.
@@ IV-B3 Analysis of results
By analyzing the experimental data of the quantized LLM on zero-shot tasks and perplexity, it can be demonstrated that evaluating the layer importance of LLMs by constructing the top-$k$ token sets corresponding to the input and output of each layer and calculating the Jaccard similarity of these sets is more effective than using cosine similarity. This evaluation approach adeptly captures the interplay between each layer and semantic information. Consequently, it facilitates a more nuanced and accurate appraisal of the significance of each layer.
This finding reveals that compared to LWQ, the LSAQ quantization method is more effective in maintaining LLM accuracy, as it loses less critical information during the quantization process. Consequently, this suggests that LSAQ has potential advantages in the field of LLM compression and deployment on edge devices.
IV-C Quantized LLM Deployment
In addition to quantization, dynamic deployment of LLMs on edge devices is also one of the core capabilities of LSAQ. To investigate whether LSAQ meets the criteria for the successful deployment of LLMs on edge devices, we performed tests on the memory requirements for loading LLM weights of the quantized Llama-2-7B/13B and Llama-3-8B. Besides, we examined the quantization strategies provided by our system for the Llama-2-7B model under various memory constraints to assess whether they are compliant with the requirements.
Figure 4 demonstrates the amount of memory required to load the weights of three LLMs at different quantization precisions. When these LLMs operate at FP16 precision, they require approximately 12.82GB, 24.36GB, and 15.14GB of memory, respectively. With the reduction of the average number of bits after quantization, the memory occupancy of all LLMs shows significant decreases. At their lowest, the memory requirements can be reduced to about 3.56GB, 6.79GB, and 5.18GB, respectively, which allows these LLMs to be easily deployed on mainstream graphics cards currently available on the market.
@@ IV-C Quantized LLM Deployment
TABLE IV: Quantization strategies for Llama-2-7B under various memory conditions.
@@
During the LLM inference process, intermediate computational results are generated, which also occupy a certain amount of memory. Therefore, when designing the quantization strategy, we reserved a portion of the storage for its inference based on the requirements of LLMs of different sizes.
Table IV illustrates the actual quantization strategies employed by LSAQ for deploying Llama-2-7B under various memory configurations. For instance, if a GPU has 16GB of available memory, it is entirely feasible to deploy Llama-2-7B directly on the GPU at FP16 precision without the need for quantization. However, when the GPU has only 6GB of available memory, it becomes necessary to quantize the least important 10 out of the 32 layers of Llama-2-7B to INT4 precision, while the remaining 22 layers are quantized to INT8 precision, thereby facilitating the deployment of the LLM.
@@ V Conclusion
This paper introduces LSAQ, a Layer-Specific Adaptive Quantization and dynamic deployment system for LLMs. LSAQ adjusts quantization strategies on hardware platforms with varying resource sizes by establishing a novel layer importance assessment mechanism, thereby significantly reducing the LLMs’ storage requirements while maintaining LLM performance as much as possible. Experimental results demonstrate that the accuracy of the LSAQ-quantized LLM on zero-shot tasks and the perplexity of the LLM are superior to existing quantization methods of the same granularity, even at lower bit widths, thus maintaining high performance. This indicates that LSAQ has potential advantages in LLM compression and on-device LLM deployment. By employing intelligent adaptive quantization methods, LSAQ can efficiently deploy quantized LLMs in resource-limited environments, providing a new perspective and approach for the deployment of LLMs on edge devices.
VI Acknowledgement
This work was supported by the National Key Research and Development Program under Grant 2024YFB4506200, the Science and Technology Innovation Program of Hunan Province under Grant 2024RC1048, and the National Key Laboratory Foundation Project under Grant 2024-KJWPDL-14.